The data set below represents bugs found by a software tester in her product during different phases of testing: 88,84,81,94,91,98,98,200. The measures of central tendency are given below: Mean:104.25, Median: 92.5,Mode98. Identify the outlier and the measurs of central tendency that is affected by the outlier.
Answer #1

Given,

Mean = 104.25,

Median = 92.5,

Mode = 98

Q1 = (84+88)/2 = 86

Q3 = (98+98)/2 = 98

Q = Q3 - Q1 = 98 - 86 = 12

Q^2 = 12 * 12 = 144

Q3 + Q^2 = 98+144 = 242

Q1 - Q^2 = 86-144 = -58

If the data beyond -58 to 242, those are the outliers. In the given data there is no outliers.

There is no data affected by the outlier. The measure of central tendency values are 81, 84, 88, 91, 94, 98, 98, 200 then mean, median, mode are calculated as

Mean = 104.25 Median = 92.5 Mode = 98.